the ques is like dis


3 raised to 2003...n dis value divided by 8...den take d fractional part {} of d ans...{3^2003 / 8}....n pls also suggest sum shortcut for dis ques if possible...

Grade:11

1 Answers

aditi jangid
13 years ago

i calculated and got 0.375 .

3^2003 = 3 (3^2002) = 3(9^1001) = 3 ( (8+1)^1001)

apply expansion.

all terms vil b divisible... only last mein 3(1001C1001) vil b left whiich is equal to 3

divide it by 8 and get ur answer....

n m sory i dnt knw ny shortcut..

hope i got the rite answer...